# Icy or Rocky? Convective or Stable?

### v1

@@
We present a new framework for constructing agnostic and yet physical models for planetary interiors and apply it to Uranus and Neptune. Unlike previous research that either impose rigid assumptions or rely on simplified empirical profiles, our approach bridges both paradigms. Starting from randomly generated density profiles, we apply an iterative algorithm that converges toward models that simultaneously satisfy hydrostatic equilibrium, match the observed gravitational moments, and remain thermodynamically and compositionally consistent. The inferred interior models for Uranus and Neptune span a wide range of possible interior structures, in particular encompassing both water-dominated and rock-dominated configurations (rock-to-water mass ratios between 0.04–3.92 for Uranus and 0.20–1.78 for Neptune). All models contain convective regions with ionic water and have temperature–pressure profiles that remain above the demixing curves for hydrogen–helium–water mixtures. This offers both a plausible explanation for the observed non-dipolar magnetic fields and indicates that no hydrogen–helium–water demixing occurs. We find a higher H-He mass fraction in the outermost convection zones for Uranus (0.62–0.73) compared to Neptune (0.25–0.49) and that Uranus’ magnetic field is likely generated deeper in the interior compared to Neptune. We infer upper limits of 0.69–0.74 (Uranus) vs. 0.78–0.92 (Neptune) for the outer edges of the dynamo regions in units of normalised radii. Overall, our findings challenge the conventional classification of Uranus and Neptune as ”ice giants” and underscore the need for improved observational data or formation constraints to break compositional degeneracy.
@@ 1 Introduction
The outer solar system harbours intriguing, and yet the least understood planets in the Solar System. Uranus and Neptune, often referred to as ”ice giants”, remain enigmatic in terms of their composition and formation and evolution history (for example Helled, 2025, and references therein). These two planets have always been key to understanding the solar system, but their importance has become even more apparent in recent years. Intermediate-mass exoplanets, with masses between Earth and Neptune, are found to be the most common planets in the galaxy (for example Zhu & Dong, 2021). A better understanding of these giants could provide answers regarding several fundamental questions such as: What mechanisms are most relevant for the formation of a planet at a given distance from a given star? Can we quantify the relation between the observable atmospheres and the hidden interiors of a planet? How does the evolution of a planet after formation influence what we can observe today? To answer such questions, advanced planet formation and evolution models are needed. Such models can be improved by obtaining constraints from planetary interior models. In particular, it is important to ensure that such interior models are built with a minimal set of assumptions. They should provide an unbiased view regarding the possible internal structures of the planets. In this work, we present a novel and agnostic method to generate self-consistent planetary interior models by improving over previous work (Morf et al., 2024).
@@
expressed by gravitational moments $J_{2n}$. We assume a north-south and azimuthally symmetric planet described by spherical coordinates $\@vec{r}=(r,\vartheta,\varphi)$. $P_{2n}$ denote the Legendre polynomials that appear during the expansion of $V$ using spherical harmonics (for example Zharkov & Trubitsyn, 1978). Note that the reference radius $R_{\text{ref}}$ is an arbitrary normalisation constant that can be equal to $R_{\text{eq}}$, but does not have to be. The gravitational moments $J_{2n}$ have only been measured for planets within our solar system, where Voyager 2 (for example French et al., 1988) provided the first data for Uranus and Neptune. These values have later been improved using moon and ring observations (Jacobson, 2009, 2014; Wang et al., 2023; French et al., 2024; Jacobson & Park, 2025). Note that in a corotating reference frame, the total exterior potential $U_{\text{ext}}$ is then given by:
@@
We briefly summarise recent progress in interior modelling of Uranus or Neptune. This includes work with so-called ”physical” and ”empirical” models. Physical models are self-consistent and provide a complete picture of a planet’s interior. They include a density, pressure, temperature, and composition profile compatible with a given Equation of State (EoS). However, physical models strongly depend on the assumptions used by the modeller. They only provide a single and rather biased possibility among many into a given planet’s interior. Empirical models a priori provide only the pressure and density profile within a planet. The interior temperature and composition can be inferred in a second step with an algorithm. Since density and pressure were first inferred without any EoS knowledge, such algorithms typically cannot adhere to the same standards as used for physical modelling. Empirical results for the temperature and composition hence suffer from inconsistencies or unphysical features (for example Neuenschwander et al., 2024; Morf et al., 2024). However, empirical models employ a minimal set of assumptions, enabling the modellers to achieve a more general view of the possible solutions for the planetary interior.
@@
Recent work on physical interior models of Uranus or Neptune includes Militzer (2024); Cano Amoros et al. (2024); Tejada Arevalo (2025). Militzer (2024) based their models on ab-initio simulations for a mixture of water, methane, and ammonia. They showed that these components phase separate under the pressure–temperature conditions in the interiors of Uranus and Neptune. The inferred interiors that match the observed gravity field and are compatible with magnetic field measurements are found to be rock-poor: 2 % and 7 % of the total mass for Uranus and Neptune, respectively. Cano Amoros et al. (2024) constructed hydrogen-water phase diagrams from available experimental and computational data. They computed interior adiabatic structure models for Uranus and Neptune and inferred a strong water depletion in the top layers of both planets. This is caused by the immiscibility of hydrogen and water at certain conditions. Tejada Arevalo (2025) presented non-adiabatic and inhomogeneous evolution models for Uranus and Neptune assuming an interior composition of hydrogen, helium, methane, ammonia, water, and rocks. By setting the relevant initial parameters, they produced a Uranus model that preserves much of its primordial internal heat. In contrast, their Neptune model undergoes adiabatic cooling of the outer envelope, in agreement with observations.
@@
Recent work on empirical interior models of Uranus or Neptune includes Movshovitz & Fortney (2022); Neuenschwander et al. (2024); Malamud et al. (2024); Morf et al. (2024); Lin et al. (2025); Mankovich et al. (2025). Movshovitz & Fortney (2022) investigated the constraining power of a high-precision orbiter measuring the gravitational fields of Uranus and Neptune. They found that while the ambiguity of Uranus’ and Neptune’s compositional profiles will still persist with high-precision data, determining the locations of any large composition gradient regions could become feasible. Neuenschwander et al. (2024) modelled the density of Uranus with up to three polytropes. Their models for Uranus’ interior include non-adiabatic regions. This leads to significantly hotter internal temperatures and higher rock-to-water ratios compared to physical models. Malamud et al. (2024) showed that Uranus and Neptune could have accreted refractory-dominated planetesimals, while still remaining icy. They investigated chemical reactions of organic-rich refractory materials and the hydrogen in gaseous atmospheres of protoplanets. Uranus and Neptune could hence consist of large amounts of methane ”ice”. Morf et al. (2024) developed a random algorithm to interpret empirical models of Uranus in terms of their temperature and composition. Additionally, they developed the Theory of Figures (Zharkov & Trubitsyn, 1978) to tenth order, allowing for accurate calculations of higher-order gravitational moments. These models in particular constrained the properties of the ionic water convection zone, likely responsible for Uranus’ magnetic field. Lin et al. (2025) argued that high degrees of mixing are required for Uranus interior models to be consistent with the measured gravitational moments. Furthermore, they showed that the gravity measurements of a future orbiter could distinguish between high and low atmospheric metallicity scenarios. Mankovich et al. (2025) calculated the spectra that could manifest in resonances with ring orbits or in Doppler imaging of Uranus’s visible surface. Their approach, originally developed for stellar oscillations, has the potential to significantly reduce the degeneracy in the set of otherwise allowed interior profiles.
@@
The aim of this study is to develop methods that generate interior models that are both agnostic and physical, to bridge the gap between physical and empirical modelling. Our paper is organized as follows: In Section 2 we introduce our methods in more detail and give a general overview of our algorithm. In Sections 3 and 4 we show the inferred interior models of Uranus and Neptune, respectively. In Section 5, we elaborate and compare the inferred properties for Uranus and Neptune such as their hydrogen-helium abundances and rock-to-water ratios (Section 5.1). We investigate the models’ consistency with the observed magnetic fields (Section 5.2) and immiscibility constraints from hydrogen-helium-water mixtures (Section 5.3). A direct comparison of the Uranus vs. Neptune models can be found in Section 5.4. We discuss the limitations of the work and the required follow-up work in Section 6. Our conclusions are presented in Section 7.
@@ 2 Methods
Interior models are designed to fit the measured planetary data. In this study we focus on Uranus and Neptune and use the data listed in Table 1.
Figure 1: An overview of our global algorithm to infer agnostic and self-consistent planetary interior models. The difference between $\tilde{\rho}$ used by the compositional algorithm and $\rho$ compatible with the interior pressure and planetary mass decreases until it vanishes using our iterative process. Traditionally, empirical models only use a single iteration, leading to inconsistencies. See Figure B.1 for more details.
We present a global algorithm to generate agnostic and physical interior models. A schematic overview of the algorithm is shown in Figure 1. A more detailed description is presented in Figure B.1. The global algorithm starts by generating an initial random density profile $\rho^{(0)}(r)$: Assume a perfectly spherical planet with radius $R$ and mass $M$. Spherical symmetry allows to describe the interior by a one-dimensional density function $\rho=\rho(r)$, where $r$ denotes the distance to the centre. Regarding the form of said function, every monotonically decreasing function $f:r\in[0,R]\rightarrow f(r)\in[1,0]$ can be rescaled to a one-dimensional density profile $\rho$ via
@@
We generate a random initial function $f$ analogous to Podolak et al. (2022): Start defining $f$ with the two tuples $(r=0,f(r)=1)$ and $(r=R,f(r)=0)$. These two tuples are interpreted as corners of the rectangle $0,R]\times[1,0]$. Now choose a random point $(r^{*},f(r^{*}))$ in the rectangle, adding a third tuple to the definition of $f$. The point $(r^{*},f(r^{*}))$ defines two new and smaller rectangles: $0,r^{*}]\times[1,f(r^{*})]$ and $r^{*},R]\times[f(r^{*}),0]$. Within these smaller rectangles, we can again choose new tuples, increasing the number of points where $f$ is defined from three to five. Now repeat this procedure by choosing new tuples for each new rectangle as often as desired. This procedure ensures that $f$ is indeed monotonically decreasing and generated in a random manner. For this work, we choose $1+2+4+8+16=31$ tuples in addition to the two initial tuples and use linear interpolation between the tuples to get to $N=512$ equidistantly spaced radial points with a corresponding value of $f$. Applying a Gaussian filter to $f$ (as in Podolak et al., 2022) and using Equation 3 yields the initial random density $\rho^{(0)}(r)$.
As a next step, $\rho^{(0)}(r)$ is given to the Theory of Figures [^1] See https://doi.org/10.5281/zenodo.16902935 for the numerical implementation. (Zharkov & Trubitsyn, 1978; Morf et al., 2024). The Theory of Figures (ToF) takes the one-dimensional density function $\rho(r)$, the mass $M$, the equatorial radius $R_{\text{eq}}$, and the rotation period $P_{\text{rot}}$ as inputs and calculates the implied two-dimensional spheroidal shape of the planet. In particular, the gravitational moments $J_{n,\text{ToF}}$ and the internal pressure $P_{\text{ToF}}(r)$ are calculated as an output. $P_{\text{ToF}}(r)$ is calculated by integrating the hydrostatic equilibrium equation:
@@
where $U$ is the total interior potential consisting of gravitational and centrifugal contributions (see appendix A.1 in Morf et al., 2024, for details).
Having obtained initial density $\rho^{(0)}(r)$ and pressure $P_{\text{ToF}}^{(0)}(r)$ profiles, we now call the compositional algorithm. The details of the compositional algorithm are discussed in Appendix B. For now, it suffices to say that the compositional algorithm uses Equations of State (EoS) for different materials to infer a temperature $T_{\text{EoS}}^{(0)}(r)$ and composition profile $\@vec{X}_{\text{EoS}}^{(0)}(r)$ based on the given density-pressure profile. However, the densities implied by the EoS $\rho_{\text{EoS}}^{(0)}(r)$ can deviate from the originally provided densities $\rho^{(0)}(r)$. Although the compositional algorithm tries to keep the deviations as small as possible, they are necessary because a random density profile will almost never represent a physically realistic interior.
Due to these deviations, we now have to renormalize the density profile (like in Equation 3) to ensure that the planet still has the correct mass $M$. We call the renormalised density profile $\rho^{(1)}(r)$. By using the ToF a second time, we get new values for the pressure $P_{\text{ToF}}^{(1)}(r)$ and gravitational moments $J^{(1)}_{n,\text{ToF}}$ that are consistent with $\rho^{(1)}(r)$. At this point, we hence have a self-consistent tuple:
@@
Figure 2: The inferred composition and convective and radiative structures for the four Uranus interior models. The legends summarise the total mass fractions for each component. The uppermost small arc of each slice is empty because we do not infer a composition for pressures below 100 bars.
However, $\rho^{(1)}(r)$ should be much more compatible with the EoS than $\rho^{(0)}(r)$ was initially. If we call the compositional algorithm again, we can expect that it can more easily infer temperature $T_{\text{EoS}}^{(1)}(r)$ and composition $\@vec{X}_{\text{EoS}}^{(1)}(r)$ profiles, whose implied densities $\rho_{\text{EoS}}^{(1)}(r)$ are close to the provided densities $\rho^{(1)}(r)$. Consequently, $\rho^{(2)}(r)$ and $\rho^{(1)}(r)$ should be much closer to each other compared to how close $\rho^{(1)}(r)$ is to $\rho^{(0)}(r)$. Therefore, by repeating the above outlined procedure, eventually there should be a $k\in\mathbb{N}$, where the difference between $\rho^{(k+1)}(r)$ and $\rho^{(k)}(r)$ is so small for all $r$, that the tuple:
@@
is practically self-consistent. Then, Equation 7 is a complete physical model of a planetary interior which has been generated in a random manner. We employ the criterion:
@@
to check whether the solution is self-consistent. Since the density and pressure are directly linked to each other via Equation 4, there is no need to define a second criterion for the density. In this work, we choose $\epsilon=0.02$. This is a conservative number given the uncertainties associated with the EoS of the different materials and their mixtures. For example, the assumption of ideal mixing (Equation 10) alone can introduce deviations of up to 5% in density (Darafeyeu et al., 2024). For a detailed discussion about algorithm convergence, see Appendix C.
Finally, note that in general $J^{(k+1)}_{n,\text{ToF}}\neq J^{(0)}_{n,\text{ToF}}$ at the end of the global algorithm. So there is no point in ensuring a precise match between the initial gravitational moments $J^{(0)}_{n,\text{ToF}}$ and the available data in Table 1. Only $J^{(k+1)}_{n,\text{ToF}}$ needs to be data consistent, promoting Equation 7 from a physical interior model to a model also compatible with the data of the considered planet.
To summarise, we iterate until the density, pressure, temperature and composition profile have converged and all agree with the data in Table 1 and the used EoS. Our method is capable of generating interior models that are both agnostic and self-consistent. Being agnostic refers to using a minimal set of assumptions like for empirical models. However, we still maintain a level of self-consistency typical for physical models. As shown below, our results highlight that no single interior model can be regarded as the standard for a given planet. A wide range of self-consistent models must be considered to get a full picture that can reliably inform formation and evolution models.

### v2

@@
We present a new framework for constructing agnostic and yet physical models for planetary interiors and apply it to Uranus and Neptune. Unlike previous research that either impose rigid assumptions or rely on simplified empirical profiles, our approach bridges both paradigms. Starting from randomly generated density profiles, we applied an iterative algorithm that converges towards models that simultaneously satisfy hydrostatic equilibrium, match the observed gravitational moments, and remain thermodynamically and compositionally consistent. The inferred interior models for Uranus and Neptune span a wide range of possible interior structures, in particular encompassing both water-dominated and rock-dominated configurations (rock-to-water mass ratios between 0.04–3.92 for Uranus and 0.20–1.78 for Neptune). All models contain convective regions with ionic water and have temperature–pressure profiles that remain above the demixing curves for hydrogen–helium–water mixtures. This offers both a plausible explanation for the observed non-dipolar magnetic fields and indicates that no hydrogen–helium–water demixing occurs. We find a higher H-He mass fraction in the outermost convection zones for Uranus (0.62–0.73) compared to Neptune (0.25–0.49) and that Uranus’ magnetic field is likely generated deeper in the interior compared to Neptune. We infer upper limits of 0.69–0.74 (Uranus) versus 0.78–0.92 (Neptune) for the outer edges of the dynamo regions in units of normalised radii. Overall, our findings challenge the conventional classification of Uranus and Neptune as ’ice giants’ and underscore the need for improved observational data or formation constraints to break compositional degeneracy.
@@ 1 Introduction
The outer solar system harbours intriguing, and yet the least understood planets in the Solar System. Uranus and Neptune, often referred to as ’ice giants’, remain enigmatic in terms of their composition and formation and evolution history (for example Helled, 2025, and references therein). These two planets have always been key to understanding the solar system, but their importance has become even more apparent in recent years. Intermediate-mass exoplanets, with masses between Earth and Neptune, are found to be the most common planets in the galaxy (for example Zhu & Dong, 2021). A better understanding of these giants could provide answers to several fundamental open problems in planetary physics. For example, it could reveal the mechanisms most relevant for the formation of a planet at a given distance from a specific star. It might help quantify the relationship between the observable atmospheres and the hidden interiors of a planet. And it could shed light on how a planet’s evolution after formation influences what we are able to observe today. To find answers to such problems, advanced planet formation and evolution models are needed. Such models can be improved by obtaining constraints from planetary interior models. In particular, it is important to ensure that such interior models are built with a minimal set of assumptions. They should provide an unbiased view regarding the possible internal structures of the planets. In this work, we present a novel and agnostic method to generate self-consistent planetary interior models by improving over previous work (Morf et al., 2024).
@@
expressed by gravitational moments $J_{2n}$. We assume a north-south and azimuthally symmetric planet described by spherical coordinates $\@vec{r}=(r,\vartheta,\varphi)$. The Legendre polynomials are denoted as $P_{2n}$ and appear during the expansion of $V$ using spherical harmonics (for example Zharkov & Trubitsyn, 1978). The reference radius $R_{\text{ref}}$ is an arbitrary normalisation constant that can be equal to $R_{\text{eq}}$, but does not have to be. The gravitational moments $J_{2n}$ have only been measured for planets within our solar system, where Voyager 2 (for example French et al., 1988) provided the first data for Uranus and Neptune. These values have later been improved using moon and ring observations (Jacobson, 2009, 2014; Wang et al., 2023; French et al., 2024; Jacobson & Park, 2025). In a corotating reference frame, the total exterior potential $U_{\text{ext}}$ is then given by:
@@
We briefly summarise recent progress in interior modelling of Uranus or Neptune. This includes work with so-called physical and empirical models. Physical models are self-consistent and provide a complete picture of a planet’s interior. They include a density, pressure, temperature, and composition profile compatible with a given Equation of State (EoS). However, physical models strongly depend on the assumptions used by the modeller. They only provide a single and rather biased possibility among many into a given planet’s interior. Empirical models a priori provide only the pressure and density profile within a planet. The interior temperature and composition can be inferred in a second step with an algorithm. Since density and pressure were first inferred without any EoS knowledge, such algorithms typically cannot adhere to the same standards as used for physical modelling. Empirical results for the temperature and composition hence suffer from inconsistencies or unphysical features (for example Neuenschwander et al., 2024; Morf et al., 2024). However, empirical models employ a minimal set of assumptions, enabling the modellers to achieve a more general view of the possible solutions for the planetary interior.
@@
Recent work on physical interior models of Uranus or Neptune includes Militzer (2024); Cano Amoros et al. (2024); Tejada Arevalo (2025). Militzer (2024) base their models on ab-initio simulations for a mixture of water, methane, and ammonia. They show that these components phase separate under the pressure–temperature conditions in the interiors of Uranus and Neptune. The inferred interiors that match the observed gravity field and are compatible with magnetic field measurements are rock-poor: 2 % and 7 % of the total mass for Uranus and Neptune, respectively. Cano Amoros et al. (2024) provide hydrogen-water phase diagrams from available experimental and computational data. They find interior adiabatic structure models for Uranus and Neptune and infer a strong water depletion in the top layers of both planets. This is caused by the immiscibility of hydrogen and water at certain conditions. Tejada Arevalo (2025) present non-adiabatic and inhomogeneous evolution models for Uranus and Neptune assuming an interior composition of hydrogen, helium, methane, ammonia, water, and rocks. By setting the relevant initial parameters, they produce a Uranus model that preserves much of its primordial internal heat. In contrast, their Neptune model undergoes adiabatic cooling of the outer envelope, in agreement with observations.
@@
Recent work on empirical interior models of Uranus or Neptune includes Movshovitz & Fortney (2022); Neuenschwander et al. (2024); Malamud et al. (2024); Morf et al. (2024); Lin et al. (2025); Mankovich et al. (2025). Movshovitz & Fortney (2022) investigate the constraining power of a high-precision orbiter measuring the gravitational fields of Uranus and Neptune. They find that while the ambiguity of Uranus’ and Neptune’s compositional profiles will still persist with high-precision data, determining the locations of any large composition gradient regions could become feasible. Neuenschwander et al. (2024) model the density of Uranus with up to three polytropes. Their models for Uranus’ interior include non-adiabatic regions. This leads to significantly hotter internal temperatures and higher rock-to-water ratios compared to physical models. Malamud et al. (2024) show that Uranus and Neptune could have accreted refractory-dominated planetesimals, while still remaining icy. They investigate chemical reactions of organic-rich refractory materials and the hydrogen in gaseous atmospheres of protoplanets. Uranus and Neptune could hence consist of large amounts of methane ’ice’. Morf et al. (2024) provide a random algorithm to interpret empirical models of Uranus in terms of their temperature and composition. Additionally, they supply the Theory of Figures (Zharkov & Trubitsyn, 1978) to tenth order, allowing for accurate calculations of higher-order gravitational moments. Their models in particular constrain the properties of the ionic water convection zone, likely responsible for Uranus’ magnetic field. Lin et al. (2025) argue that high degrees of mixing are required for Uranus interior models to be consistent with the measured gravitational moments. Furthermore, they show that the gravity measurements of a future orbiter could distinguish between high and low atmospheric metallicity scenarios. Mankovich et al. (2025) provide the spectra that could manifest in resonances with ring orbits or in Doppler imaging of Uranus’s visible surface. Their approach, originally developed for stellar oscillations, has the potential to significantly reduce the degeneracy in the set of otherwise allowed interior profiles.
@@
The aim of this study is to develop methods that generate interior models that are both agnostic and physical, to bridge the gap between physical and empirical modelling. Our paper is organised as follows: In Section 2 we introduce our methods in more detail and give a general overview of our algorithm. In Sections 3 and 4 we show the inferred interior models of Uranus and Neptune, respectively. In Section 5, we elaborate and compare the inferred properties for Uranus and Neptune such as their hydrogen-helium abundances and rock-to-water ratios (Section 5.1). We investigate the models’ consistency with the observed magnetic fields (Section 5.2) and immiscibility constraints from hydrogen-helium-water mixtures (Section 5.3). A direct comparison of the Uranus versus Neptune models can be found in Section 5.4. We discuss limitations and the required follow-up work in Section 6 and present conclusions in Section 7.
@@ 2 Methods
Interior models are designed to fit the measured planetary data. In this study we focused on Uranus and Neptune and used the data listed in Table 1.
Figure 1: Overview of our global algorithm to infer agnostic and self-consistent planetary interior models. The difference between $\tilde{\rho}$ used by the compositional algorithm and $\rho$ compatible with the interior pressure and planetary mass decreases until it vanishes using our iterative process. Traditionally, empirical models only use a single iteration, leading to inconsistencies. Figure B.1 contains more details.
We present a global algorithm to generate agnostic and physical interior models. A schematic overview of the algorithm is shown in Figure 1. A more detailed description is presented in Figure B.1. The global algorithm starts by generating an initial random density profile $\rho^{(0)}(r)$: We assume a perfectly spherical planet with radius $R$ and mass $M$. Spherical symmetry allows to describe the interior by a one-dimensional density function $\rho=\rho(r)$, where $r$ denotes the distance to the centre. Regarding the form of said function, every monotonically decreasing function $f:r\in[0,R]\rightarrow f(r)\in[1,0]$ can be rescaled to a one-dimensional density profile $\rho$ via
@@
We generated a random initial function $f$ analogous to Podolak et al. (2022): Start defining $f$ with the two tuples $(r=0,f(r)=1)$ and $(r=R,f(r)=0)$. These two tuples are interpreted as corners of the rectangle $0,R]\times[1,0]$. Now choose a random point $(r^{*},f(r^{*}))$ in the rectangle, adding a third tuple to the definition of $f$. The point $(r^{*},f(r^{*}))$ defines two new and smaller rectangles: $0,r^{*}]\times[1,f(r^{*})]$ and $r^{*},R]\times[f(r^{*}),0]$. Within these smaller rectangles, we can again choose new tuples, increasing the number of points where $f$ is defined from three to five. Now repeat this procedure by choosing new tuples for each new rectangle as often as desired. This procedure ensures that $f$ is indeed monotonically decreasing and generated in a random manner. For this work, we chose $1+2+4+8+16=31$ tuples in addition to the two initial tuples and used linear interpolation between the tuples to get to $N=512$ equidistantly spaced radial points with a corresponding value of $f$. Applying a Gaussian filter to $f$ (as in Podolak et al., 2022) and using Equation 3 yielded the initial random density $\rho^{(0)}(r)$.
As a next step, $\rho^{(0)}(r)$ was given to the Theory of Figures [^1] https://doi.org/10.5281/zenodo.16902935 provides the numerical implementation. (Zharkov & Trubitsyn, 1978; Morf et al., 2024). The Theory of Figures (ToF) takes the one-dimensional density function $\rho(r)$, the mass $M$, the equatorial radius $R_{\text{eq}}$, and the rotation period $P_{\text{rot}}$ as inputs and calculates the implied two-dimensional spheroidal shape of the planet. In particular, the gravitational moments $J_{n,\text{ToF}}$ and the internal pressure $P_{\text{ToF}}(r)$ are calculated as an output. The pressure $P_{\text{ToF}}(r)$ is calculated by integrating the hydrostatic equilibrium equation:
@@
where $U$ is the total interior potential consisting of gravitational and centrifugal contributions (Appendix A.1 in Morf et al., 2024, for details).
Having obtained initial density $\rho^{(0)}(r)$ and pressure $P_{\text{ToF}}^{(0)}(r)$ profiles, the compositional algorithm is called next. The details of the compositional algorithm are discussed in Appendix B. In a nutshell, the compositional algorithm uses Equations of State (EoS) for different materials to infer a temperature $T_{\text{EoS}}^{(0)}(r)$ and composition profile $\@vec{X}_{\text{EoS}}^{(0)}(r)$ based on the given density-pressure profile. However, the densities implied by the EoS $\rho_{\text{EoS}}^{(0)}(r)$ can deviate from the originally provided densities $\rho^{(0)}(r)$. Although the compositional algorithm tries to keep the deviations as small as possible, they are necessary because a random density profile will almost never represent a physically realistic interior.
Due to these deviations, the density profile has to be renormalised (like in Equation 3) to ensure that the planet still has the correct mass $M$. The renormalised density profile is denoted as $\rho^{(1)}(r)$. Using the ToF a second time provides new values for the pressure $P_{\text{ToF}}^{(1)}(r)$ and gravitational moments $J^{(1)}_{n,\text{ToF}}$ that are consistent with $\rho^{(1)}(r)$. At this point, one hence has a self-consistent tuple:
@@
Figure 2: Inferred composition and convective and radiative structures for the four Uranus interior models. The legends summarise the total mass fractions for each component. The uppermost small arc of each slice is empty because we do not infer a composition for pressures below 100 bars.
However, $\rho^{(1)}(r)$ should be much more compatible with the EoS than $\rho^{(0)}(r)$ was initially. By calling the compositional algorithm again, one can expect that it can more easily infer temperature $T_{\text{EoS}}^{(1)}(r)$ and composition $\@vec{X}_{\text{EoS}}^{(1)}(r)$ profiles, whose implied densities $\rho_{\text{EoS}}^{(1)}(r)$ are close to the provided densities $\rho^{(1)}(r)$. Consequently, $\rho^{(2)}(r)$ and $\rho^{(1)}(r)$ should be much closer to each other compared to how close $\rho^{(1)}(r)$ is to $\rho^{(0)}(r)$. Therefore, by repeating the above outlined procedure, eventually there should be a $k\in\mathbb{N}$, where the difference between $\rho^{(k+1)}(r)$ and $\rho^{(k)}(r)$ is so small for all $r$, that the tuple:
@@
is practically self-consistent. Then, Equation 7 is a complete physical model of a planetary interior which has been generated in a random manner. We employed the criterion:
@@
to check whether a solution is self-consistent. Since the density and pressure are directly linked to each other via Equation 4, there is no need to define a second criterion for the density. In this work, we chose $\epsilon=0.02$. This is a conservative number given the uncertainties associated with the EoS of the different materials and their mixtures. For example, the assumption of ideal mixing (Equation 10) alone can introduce deviations of up to 5% in density (Darafeyeu et al., 2024). A detailed discussion about algorithm convergence is given in Appendix C.
Finally, in general $J^{(k+1)}_{n,\text{ToF}}\neq J^{(0)}_{n,\text{ToF}}$ holds true at the end of the global algorithm. So there is no point in ensuring a precise match between the initial gravitational moments $J^{(0)}_{n,\text{ToF}}$ and the available data in Table 1. Only $J^{(k+1)}_{n,\text{ToF}}$ has to be data consistent, promoting Equation 7 from a physical interior model to a model also compatible with the data of the considered planet.
To summarise, we iterated until the density, pressure, temperature, and composition profile have converged and all agreed with the data in Table 1 and the used EoS. Our method is capable of generating interior models that are both agnostic and self-consistent. Being agnostic refers to using a minimal set of assumptions like for empirical models. However, we still maintain a level of self-consistency typical for physical models. As shown below, our results highlight that no single interior model can be regarded as the standard for a given planet. A wide range of self-consistent models must be considered to get a full picture that can reliably inform formation and evolution models.